### v1

# Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis

Abstract.
Prompt optimization has emerged as an effective alternative to retraining for improving the performance of Large Language Models (LLMs). However, most existing approaches treat evaluation as a black box, relying solely on numerical scores while offering limited insight into why a prompt succeeds or fails. They also depend heavily on trial-and-error refinements, which are difficult to interpret and control. In this paper, we introduce MA-SAPO, a Multi-Agent framework for Score-Aware Prompt Optimization. Compared to prior methods, MA-SAPO explicitly couples evaluation outcomes with structured reasoning to guide systematic edits. The framework specifically consists of two stages: during the Reasoning Phase, agents collaboratively explain metric scores, diagnose weaknesses, and synthesize targeted refinements that are stored as reusable reasoning assets; during the Test Phase, agents retrieve these assets to analyze optimized prompts and apply only evidence-grounded edits. By turning evaluation signals into interpretable reasoning chains, MA-SAPO produces prompt refinements that are more transparent, auditable, and controllable. Experiments on the HelpSteer1/2 benchmarks demonstrate consistent improvements over single-pass prompting, retrieval-augmented baselines, and prior multi-agent strategies, validating the effectiveness of our approach. [^1] Code is available at: https://anonymous.4open.science/r/MA-SAPO-F313.
Automatic Prompt Optimization, Score-Aware Prompt Optimization, Retrieval-Augmented Generation, Large Language Models (LLMs), Multi-Agent System
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Multi-agent systemsccs: Information systems Query representationccs: Computing methodologies Natural language generationccs: Information systems Relevance assessment
Figure 1. Comparative overview of prompt optimization methods. MA-SAPO enhances interpretability by linking evaluation scores to reasoning-driven and evidence-based refinements.
1. Introduction
Large Language Models (LLMs) have emerged as powerful tools capable of tackling a wide range of tasks, from reasoning and summarization to complex dialogue and code generation (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). However, their performance remains highly sensitive to the wording, structure, and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Errica et al., 2024; Cao et al., 2024; Razavi et al., 2025). Consequently, prompt optimization has rapidly emerged as a practical alternative for improving model behavior, as it avoids the need for costly retraining or parameter updates.
Early research explored single-pass prompt optimization strategies that treated the model itself as the optimizer. For instance, methods such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role assignment through carefully designed system instructions (Wu et al., 2023; Kong et al., 2023), or structured variants such as Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2023) demonstrated that reasoning patterns or role signals embedded into prompts could substantially improve downstream performance. While effective, these approaches typically operate in isolation: a single model instance refines prompts without incorporating systematic feedback from prior attempts or external resources.
As LLMs have advanced, multi-agent frameworks have gained traction as a way to diversify perspectives and coordinate multiple roles in the optimization process (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025). By assigning diverse specialized agents to different tasks such as decomposition, evaluation, or refinement, these systems move beyond single-pass rewriting to more structured optimization pipelines. At the same time, optimization-theoretic methods (Li et al., 2025b; Cui et al., 2025) and label-efficient pipelines (Xiang et al., 2025; Dong et al., 2025) have sought to reduce the reliance on costly feedback while ensuring scalability across tasks.
Despite these advances, several common limitations remain: (1) most existing frameworks reduce evaluation to outcome scores, treating it as a black box and leaving practitioners uncertain about why certain prompts succeed or fail; (2) optimization often relies on repetitive trial-and-error refinements that consume significant computation while offering little transparency; (3) even when reasoning is incorporated, it usually remains implicit and is not transformed into explicit, auditable artifacts that can guide systematic edits; (4) current pipelines largely optimize for higher outcome scores but offer limited interpretability and controllability, preventing users from understanding trade-offs or applying targeted adjustments.
To address these limitations, we propose MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly links metric outcomes to actionable edits through structured reasoning artifacts. Specifically, MA-SAPO operates in two complementary phases. In the reasoning phase, three agents transform annotated scores into progressively enriched artifacts: (1) a Metric Explainer Agent interprets evaluation dimensions, (2) a Diagnostician Agent uncovers error sources and trade-offs, and (3) an Action Synthesizer Agent produces concrete edit directives. These artifacts are stored as reusable reasoning assets. In the test phase, a retrieval-augmented pipeline leverages these assets: (1) an Analyzer Agent contrasts the current prompt with retrieved exemplars to identify improvement opportunities, while (2) a Refiner Agent applies targeted edits supported by diagnostic evidence, ensuring that optimizations are interpretable and controllable.
Extensive experiments on the open-source HelpSteer1/2 benchmarks (Wang et al., 2023, 2024a) demonstrate that MA-SAPO consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, highlighting the importance of structured reasoning as a bridge between evaluation and optimization. Our main contributions include:
  • We propose a score-aware multi-agent training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets, enabling transparent prompt explanations, fine-grained diagnoses, and actionable edit directives.
  • We design MA-SAPO, a modular and computationally efficient framework that achieves interpretable, auditable, and controllable improvements across diverse datasets and model backbones, while substantially reducing token and API call budgets.
  • We conduct extensive experiments on HelpSteer1 and HelpSteer2 benchmarks, demonstrating consistent gains across multiple evaluation metrics and outperforming six state-of-the-art baselines in both effectiveness and efficiency.
2. Related Work
Single-Pass Prompt Optimization. Prompt optimization has rapidly evolved as a practical means of improving the performance of LLMs without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025a). Early studies highlighted the centrality of prompts as control variables for steering model behavior. Specifically, few-shot prompting (Brown et al., 2020) demonstrated that models could adapt to unseen tasks using in-context examples alone, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) injected external knowledge to improve factuality and adaptability. These two paradigms, context construction and evidence injection, laid the foundation for single-pass prompt optimization methods. Subsequent research explored reasoning-augmented prompting. Chain-of-Thought (CoT) (Wei et al., 2022) introduced intermediate reasoning steps. Building on this idea, Tree-of-Thought (ToT) (Yao et al., 2023) generalized linear reasoning chains into branching structures with self-evaluation via lookahead and backtracking. Other structured strategies, such as Skeleton-of-Thought (SoT) (Ning et al., 2023), separate planning from execution to parallelize generation, while Graph-of-Thought (GoT) (Besta et al., 2024) organizes reasoning trajectories into graph topologies. Evolutionary and heuristic single-pass approaches have also been explored. For instance, PromptBreeder (Fernando et al., 2023) evolves a population of prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines prompts via feedback-driven critique-and-synthesis loops.
Multi-Agent Based Prompt Optimization. Beyond single-pass methods, multi-agent systems have emerged as a powerful paradigm for prompt optimization (Du et al., 2023; Liu et al., 2024; Chen et al., 2024; Li et al., 2024). Early agentic frameworks such as Self-Ask (Press et al., 2022) introduced decomposition by generating sub-questions and integrating external retrieval, laying the groundwork for cooperative prompting strategies. Also, Multi-Agent Debate (MAD) (Du et al., 2023) advanced this idea by allowing agents to iteratively present and critique answers in a roundtable format, thereby improving factual grounding and reasoning quality. Reflection-based frameworks (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by storing feedback as memory for subsequent optimization attempts. Building on this foundation, recent systems formalized agent roles and workflows to optimize prompts more systematically (Li et al., 2023). Subsequent work explicitly targeted prompt optimization as the objective. For instance, MASS (Zhou et al., 2025) searches over agent topologies while performing block-level prompt refinement, followed by global optimization. Additionally, MARS (Zhang et al., 2025) leverages a Planner and Teacher–Critic–Student dialogue to iteratively propose, critique, and revise prompts in structured rounds, and also MAPGD (Han et al., 2025) integrates pseudo-gradient feedback with bandit-style exploration, offering sample-efficient optimization through agent collaboration. However, most existing frameworks treat evaluation as a black box with limited interpretability and rely on costly trial-and-error refinements. MA-SAPO addresses this by grounding optimization in score-aware reasoning assets that link evaluation outcomes to actionable, interpretable, and auditable refinements.
Figure 2. Overview of the MA-SAPO framework. The training phase constructs reasoning assets $(\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}))$ from annotated prompt–response pairs, while the test phase retrieves top-k training examples and their reasoning assets to refine a new test prompt.
3. Methodology
3.1. Preliminaries
We denote the training dataset as
$\mathcal{D}_{train}=\{\tau_{i}\}_{i=1}^{N},\quad\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i}),$
where each entry $\tau_{i}$ consists of a prompt $p_{i}$, its associated response $r_{i}$, and a set of scores $\mathcal{S}_{i}=\{s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\}$, corresponding to the dimensions of helpfulness, correctness, coherence, complexity, and verbosity. The test set is denoted as
$\mathcal{D}_{test}=\{p_{j}\}_{j=1}^{M},$
where only prompts are available, and corresponding responses are generated during evaluation. Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can later be retrieved and utilized to optimize new prompts and their responses. Thus, the training phase serves as data generation, while the test phase operates as a retrieval-augmented optimization pipeline.
Throughout this framework, we employ several specialized agents denoted as $G$. Each agent is designed with a distinct functional role: $G_{\text{exp}}$ explains metric-level outcomes, $G_{\text{diag}}$ diagnoses weaknesses and trade-offs, $G_{\text{syn}}$ synthesizes actionable directives, $G_{\text{ana}}$ analyzes retrieved examples, and $G_{\text{ref}}$ refines prompts into optimized versions. These agents are executed in sequential order within their respective phases, and their outputs are explicitly stored for later retrieval and reasoning. The overview of the MA-SAPO framework is shown in Figure 2.
3.2. Training Phase: Reasoning Asset Construction
In the training phase, three agents are executed sequentially for each $\tau_{i}\in\mathcal{D}_{train}$. All three agents operate on the same triplet $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$, ensuring that their reasoning is consistently grounded in both the input and output of the original instance.
3.2.1. Metric Explainer Agent $G_{\text{exp}}$
The Metric Explainer Agent provides natural language justifications for the given scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
(1) $\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$
where $P_{\text{exp}}$ is the system prompt guiding explanation. Each card explicitly outlines the reason why $\tau_{i}$ received its annotated scores. The reasoning card $\mathcal{C}_{i}$ then serves as input to the next agent, together with $\tau_{i}$.
3.2.2. Diagnostician Agent $G_{\text{diag}}$
The Diagnostician Agent operates on $\tau_{i}$ and the reasoning card $\mathcal{C}_{i}$. It extends $\mathcal{C}_{i}$ by analyzing metric-level weaknesses and trade-offs. $G_{\text{diag}}$ produces a diagnostic summary $\mathcal{D}_{i}$ that identifies: (1) the key causes of low-scoring dimensions, and (2) trade-offs across metrics (e.g., verbosity vs. coherence).
(2) $\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$
The diagnostic summary $\mathcal{D}_{i}$ is then passed to the next stage, along with $\tau_{i}$.
3.2.3. Action Synthesizer Agent $G_{\text{syn}}$
The Action Synthesizer Agent receives $\tau_{i}$ together with $\mathcal{C}_{i}$ and $\mathcal{D}_{i}$. It converts these enriched insights into actionable edit directives (EDs). Each directive corresponds to a concrete modification strategy for improving $\tau_{i}$:
(3) $\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$
where $\mathcal{E}_{i}=\{e_{1},e_{2},\dots,e_{m}\}$ is a set of recommended edits.
Definition (Reasoning Assets).
For notational convenience, we define the collection of training outputs for instance $i$ as:
$\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}).$
As a result, $\mathcal{R}_{i}$ is stored as the reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase. Importantly, these reasoning assets are designed as semi-structured text, which makes them machine-parseable and directly usable by downstream agents in the test phase.
3.3. Test Phase: Retrieval and Reformulation
In the test phase, optimization is performed via retrieval-augmented generation. Given a new prompt $p_{\text{test}}$, we adopt a sparse lexical retriever that ranks training prompts based on term-frequency and inverse document-frequency statistics. Each training prompt $p_{j}$ is represented in the sparse lexical space, and a relevance score is computed between $p_{\text{test}}$ and $p_{j}$ according to token overlap weighted by discriminative importance. The top-$k$ most relevant prompt-response pairs are then retrieved along with their corresponding reasoning assets:
$\{(\tau_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
3.3.1. Analyzer Agent $G_{\text{ana}}$
The Analyzer Agent compares $p_{\text{test}}$ against retrieved examples to identify improvement points. It outputs an improvement report $\mathcal{A}$:
(4) $\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{ana}).$
The Analyzer Agent transforms raw retrieval into structured insights rather than merely transferring edits from similar prompts. The agent highlights concrete weaknesses and improvement opportunities by contrasting the test prompt with retrieved reasoning assets.
3.3.2. Refiner Agent $G_{\text{ref}}$
The Refiner Agent regenerates an optimized prompt $\hat{p}$ by incorporating the improvement report:
(5) $\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$
It operationalizes the analyzer’s insights into a concrete optimized prompt. Compared to a direct rewriting approach, which may introduce irrelevant or unjustified changes, the refiner explicitly conditions on the improvement report, producing refinements that are focused, justifiable, and consistent with diagnostic evidence. The final optimized prompt $\hat{p}$ is then used to generate a corresponding optimized response $\hat{r}$.
3.4. Evaluation
[…]
(6) $\text{Score}(\hat{p},\hat{r})=\langle s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\rangle,$
where each $s_{m}\in[0,4]$ for $m\in\mathcal{M}$. These scores are normalized into the $[0,1]$ range and then aggregated into a single composite value. [^2] This evaluation methodology is implemented in our setup in Section 4.2. Additionally, We complement automatic metrics with a human evaluation (Section 5.3) that tests (H1) the usefulness, accuracy, and consistency of the reasoning and (H2) whether optimized prompts preserve the original intent; 14 expert annotators, all data scientists or AI engineers, participated.
4. Experimental Design and Setup
4.1. Datasets
We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2023, 2024a), which provide human-annotated prompt-response pairs for evaluating prompt optimization in Large Language Models (LLMs). Both datasets share the same annotation schema across five quality dimensions, but differ in size and design.
We specifically build the retrieval corpus from the HelpSteer2 training dataset, where each instance $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$ is augmented with the reasoning assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$ produced during our training pipeline. Evaluation is conducted on both HelpSteer1 and HelpSteer2 validation prompts.
HelpSteer1 contains 35.3k training and .79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0-4 scale. It served as the first large-scale benchmark for aligning model outputs with human feedback and provides broad coverage of prompt-response styles.
HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024a). We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field.
Table 1. Overview of HelpSteer1/2 Annotation Dimensions. Each scored on a 0–4 scale.
Attribute Description
Helpfulness Overall helpfulness of the response to the prompt.
Correctness Inclusion of all pertinent facts without errors.
Coherence Consistency and clarity of expression.
Complexity Intellectual depth required to write the response (e.g., basic competency vs. domain expertise).
Verbosity Amount of detail included in the response relative to what is asked.
Reasoning–asset generation (corpus construction).
To construct the retrieval corpus, we employ the three reasoning agents ($G_{\text{exp}}$, $G_{\text{diag}}$, $G_{\text{syn}}$) using the o4-mini reasoning model from OpenAI (Achiam et al., 2023). This model is used exclusively for reasoning asset construction rather than downstream inference. All generations follow default API hyperparameters. For each training instance $\tau_{i}$, it produces a reasoning card $\mathcal{C}_{i}$, a diagnostic summary $\mathcal{D}_{i}$, and a set of edit directives $\mathcal{E}_{i}$, which together form the stored reasoning assets $\mathcal{R}_{i}$.
Generation models for evaluation.
For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model LLaMA-3-8B-Instruct (Dubey et al., 2024) as the backbone models. These models are responsible for generating (i) optimized prompts $\hat{p}$ from the Refiner Agent and (ii) the corresponding optimized responses $\hat{r}$ conditioned on $\hat{p}$. Both models are configured with temperature =0=0 to ensure deterministic decoding, and all other hyperparameters remain at their defaults.
Retrieval.
We adopt a sparse retriever based on BM25 (Robertson et al., 2009) over prompt text. For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^3] A detailed analysis of $k$ values and their effects on performance is in Section 5.1.1. and attach their paired responses and reasoning assets:
$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
4.2. Evaluation Method and Metrics
Judge model.
We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024b) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality. For any prompt-response pair, the judge model returns five HelpSteer-aligned scores (shown in Table 1):
$\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\},\quad s_{m}\in[0,4].$
Evaluation Metrics.
Each candidate pair $(p,r)$ is evaluated across the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$. We normalize scores by dividing by $4$, yielding $\tilde{s}_{m}=s_{m}/4\in[0,1]$. The overall quality score is then defined as the average over all metrics:
$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$
This produces a single composite value in the range $[0,1]$, reflecting the overall response quality.
4.3. Prompt Design
To guide each agent in MA-SAPO, we carefully designed specialized prompts tailored to their roles in both the training and test phases. [^4] For brevity, we show shortened prompts; full templates are in our code repository.
  1. (1)
    The training prompts (Metric Explainer, Diagnostician, and Action Synthesizer) focus on generating structured reasoning assets by interpreting evaluation scores, diagnosing weaknesses, and synthesizing actionable improvements.
  2. (2)
    The test prompts (Analyzer and Refiner) leverage these assets to contrast retrieved exemplars with the original prompt and iteratively refine it while preserving the user’s intent.
These prompts form the backbone of MA-SAPO’s reasoning-driven optimization pipeline, ensuring interpretability, controllability, and consistency across all stages. The prompt templates are provided below:
Prompt Templates for Different Agents in Training and Testing Phases Metric Explainer Agent Prompt You are an evaluation explainer. Given a user prompt, a model response, and five 0–4 scores (helpfulness, correctness, coherence, complexity, verbosity), write ONE cohesive paragraph (6–9 sentences) that explains WHY each score was assigned and HOW to improve the response. Rules: - Output must be a single paragraph in plain text (no lists, no JSON, no headings) (…) Diagnostician Agent Prompt You are a precise evaluation doctor for LLM outputs. Task: Write ONE cohesive paragraph that (1) diagnoses the main Strength and Weaknesses, (2) explains why they matter in terms of the five metrics, and (3) prescribes concrete fixes. Rules: - Evidence-first: rely only on the prompt + response + reasoning. - Mention each metric once (helpfulness, correctness, coherence, complexity, verbosity) with a brief, specific rationale (…) Output: One paragraph. Action Synthesizer Agent Prompt Mission: Based on the structured diagnosis (Diagnostician Agent), provide a single actionable prompt suggestion that guides improvements while preserving the original intent. Focus on highlighting strengths, addressing weaknesses, and suggesting … Task: Using the diagnosis, propose concise guidance that: 1. Maintains the core topic and original goal of the prompt. 2. Suggests ways to strengthen depth, clarity, and structure. 3. Recommends addition or adjustment (…) Analyzer Agent Prompt You are the Analyzer Agent. Task: (1) Independently evaluate the original prompt (strengths and weaknesses). (2) Analyze retrieved prompt–response–feature triples from three agents. (3) Integrate insights into a unified analysis. Rules: - Begin with direct evaluation of the original prompt. - Evaluate each agent’s output concisely (strengths/weaknesses) (…) Output: A structured but cohesive paragraph-length analysis. Refiner Agent Prompt You are the Refiner Agent. Task: Using Analyzer feedback, optimize the original prompt to resolve all issues while preserving intent. Rules: - Extract the core improvement points from the Analyzer. - Prioritize Action Synthesizer suggestions over others (…) Output: Final optimized prompt.
4.4. Baselines
We implement six baselines under three different categories: single-pass without retrieval (1, 2, and 3), retrieval-augmented without reasoning assets (4), and multi-agent prompt optimization (5 and 6). All methods generate an optimized prompt from the initial prompt, and the final response is produced from that optimized prompt by the same backbone LLM used for the method. Specifically, the six baseline methods are detailed below:
  1. (1)
    Direct Generation: The LLM rewrites the input prompt into an optimized one, without intermediate reasoning.
  2. (2)
    Chain-of-Thought (CoT) [2022] (Wei et al., 2022): The LLM performs step-by-step reasoning about how to improve the prompt, then outputs an optimized version.
  3. (3)
    Role Assignment [2023] (Shanahan et al., 2023): A crafted system role (e.g., “You are a meticulous assistant optimizing prompts for helpfulness, correctness, coherence, complexity, and verbosity.”) guides the LLM to output an optimized prompt in one pass.
  4. (4)
    Retrieval-Augmented Generation (RAG) [2020] (Lewis et al., 2020): M25 retrieves the top-$k=10$ training prompts with their responses and corresponding scores. The LLM references these exemplars to produce an optimized prompt.
  5. (5)
    MAD [2023] (Du et al., 2023): MAD (Multi-Agent Debate) framework organizes several LLM “debaters” that each propose an answer with a rationale, then engage in multi-round critiques of one another’s reasoning. A separate judge (or a consensus rule) evaluates the arguments at each round and selects the final answer, which measurably improves factuality and step-by-step reasoning on strategic-reasoning tasks.
  6. (6)
    MARS [2025] (Zhang et al., 2025): MARS is a hierarchical system of seven agents. A Planner designs the optimization trajectory, while a Teacher, Critic, Student Socratic dialogue iteratively refines the prompt. At each iteration, a Target module evaluates the candidates, logs the history, and the system finally outputs the best prompt.
Table 2. Main results on HelpSteer1 and HelpSteer2. Columns report the five HelpSteer metrics (Help, Corr, Coh, Comp, Verb) normalized to $[0,1]$ and their mean (Avg). Methods are grouped into (i) single-pass prompting, (ii) retrieval-augmented generation (no reasoning assets), and (iii) multi-agent frameworks. Best within each backbone is bold; second best is underlined.
Dataset Methods GPT-4o LLaMA-3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1 Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
Chain-of-Thought (CoT) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
Role Assignment 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
RAG (sparse, k=10k{=}10) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
MAD 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
MARS 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
MA-SAPO (Ours) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2 Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
Chain-of-Thought (CoT) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
Role Assignment 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
RAG (sparse, k=10k{=}10) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
MAD 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
MARS 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
MA-SAPO (Ours) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
For all baselines, we adopt GPT-4o and LLaMA-3-8B-Instruct as the backbone models, ensuring consistency and fairness in comparison by aligning with our main experimental setup. Baseline experiments were conducted with the following settings for fairness and fairness of our framework: all LLM temperatures were set to 0 and all other settings were set to LLM defaults. For MAD, it was fixed as $agent=2$ and $round=3$, and MARS designed the planner’s steps in steps 343-4 and set the entire iteration to $5$.
Table 3. Ablation Studies on Retrieval Depth and Test Agent Combination. Varying the number of retrieved training examples ($k$) shows that $k=3$ (our default, denoted as MA-SAPO (k=3)) provides the best balance between contextual coverage and robustness. Collapsing the Analyzer and Refiner into a single agent yields competitive but less consistent results. Best is highlighted in bold; second best is underlined.
Dataset Methods GPT-4o LLaMA-3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1 $k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
$k=2$ 0.5211 0.6287 0.8591 0.4960 0.7318 0.6473 0.4002 0.4747 0.8236 0.4726 0.8384 0.6019
$k=4$ 0.5204 0.6266 0.8606 0.4965 0.7369 0.6482 0.4028 0.4778 0.8272 0.4708 0.8389 0.6035
Test Agents Combination 0.4672 0.5415 0.8619 0.3670 0.6403 0.5756 0.4649 0.5203 0.8323 0.4363 0.7249 0.5957
MA-SAPO (k=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2 $k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
$k=2$ 0.5107 0.6044 0.8504 0.5158 0.7450 0.6452 0.3980 0.4713 0.8268 0.4845 0.8396 0.6041
$k=4$ 0.5066 0.6012 0.8480 0.5202 0.7451 0.6442 0.3947 0.4685 0.8251 0.4786 0.8395 0.6013
Test Agents Combination 0.5404 0.6012 0.8891 0.4455 0.7398 0.6432 0.4379 0.4720 0.7968 0.4918 0.7321 0.5861
MA-SAPO (k=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
5. Experimental Results and Analysis
5.1. Main Results
In our main experiments, we compare MA-SAPO with three families of baselines: (1) single-pass prompting methods including Direct Generation, Chain-of-Thought (CoT), and Role Assignment, (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks such as MAD and MARS. As outlined in Section 4.2, we evaluate on both HelpSteer1 and HelpSteer2 using the five annotated quality dimensions and their normalized average score as metrics. The results are shown in Table 2.
In detail, Single-pass methods provide the simplest form of optimization, but their weakness lies in performing edits in isolation. Direct Generation and CoT are essentially one-shot rewrites that lack diagnostic grounding, which makes their improvements brittle and often inconsistent across metrics. Role Assignment can guide the model to consider multiple aspects explicitly, yet it still suffers from over-reliance on handcrafted role descriptions and cannot adapt when those roles fail to cover task-specific subtleties. RAG improves upon these by injecting exemplars, but the optimization process remains unguided: retrieved examples are consumed without explanation of why they are useful, leaving the refinement process heuristic and prone to inconsistency. As a result, these methods either underperform on complex prompts or achieve higher scores only at the cost of verbosity and drift in task semantics.
Recent multi-agent frameworks such as MAD and MARS perform better by coordinating multiple roles in a structured workflow, yet they remain outcome-driven and heavy in cost [^5] A detailed cost and latency comparisons are provided in Section 5.4.. Their reliance on repeated debate or critique cycles incurs significant computational overhead, while still treating evaluation as a scalar target rather than a source of actionable reasoning. In contrast, our MA-SAPO consistently surpasses all baselines across datasets and backbones by explicitly converting evaluation signals into reusable reasoning assets. The Analyzer to Refiner pipeline applies only evidence-backed edits drawn from these assets, ensuring that optimizations are not only effective but also interpretable and controllable [^6] A detailed case study is provided in Section 5.5.. For instance, on HelpSteer1 with GPT-4o, MA-SAPO achieves 0.6486 average score compared to 0.5096 for MARS and 0.4784 for RAG. Similar margins hold across HelpSteer2 and with LLaMA-3-8B, confirming that structured reasoning, rather than trial-and-error search, is the main driver of performance. Overall, MA-SAPO combines accuracy and efficiency, making it a more practical and principled framework for prompt optimization than competitive baselines.
5.2. Ablation Study
5.2.1. Top-k Retrieval Variations
To examine how the number of retrieved training examples influences the performance of MA-SAPO, we vary the retrieval depth $k\in\{1,2,3,4\}$. Table 3 reports results for both GPT-4o and LLaMA-3-8B models across HelpSteer1 and HelpSteer2.
For both backbones, increasing $k$ from 1 to 3 generally improves overall performance. With $k=1$, the Analyzer is under-informed, as it can only compare the test prompt against a single exemplar, limiting the diversity of reasoning assets available. Moving to $k=2$ provides broader evidence and modest gains, while $k=3$ yields the best balance across all five metrics and achieves the highest average performance. At $k=4$, results remain competitive but show slight declines in some dimensions, suggesting that introducing too many exemplars adds noise and creates conflicting diagnostic signals.
These findings highlight a clear trade-off: smaller $k$ restricts contextual diversity, while larger $k$ risks diluting the relevance of retrieved assets. Overall, $k=3$ achieves the most consistent and robust performance across datasets and backbones, and we adopt it as the default retrieval setting in all subsequent experiments.
5.2.2. Test Agents Combination
To further assess the robustness of MA-SAPO, we conduct an ablation study where the two test-phase agents are merged into a single combined agent, referred to as Test Agents Combination. Table 3 summarizes the results for both GPT-4o and LLaMA-3-8B on HelpSteer1 and HelpSteer2.
We observe that the combined variant achieves competitive performance, occasionally surpassing the full MA-SAPO in a single metric. However, the average performance consistently falls behind the full framework. This suggests that while a unified agent can capture certain improvements, it lacks the structured diagnostic pipeline of MA-SAPO, where the Analyzer first produces evidence-grounded reports and the Refiner then executes targeted edits.
These findings demonstrate the necessity of separating the two roles in the test phase. The Analyzer to Refiner pipeline enforces a more interpretable and controllable workflow, ensuring that refinements are justified by retrieved reasoning assets rather than introduced implicitly by a single agent. As a result, the modular design of MA-SAPO is not only more reliable but also essential for achieving consistent gains across datasets and backbones.
5.3. Qualitative Analysis
To complement the automatic benchmark results, we conducted a human evaluation designed to assess both the interpretability and controllability of MA-SAPO’s optimization process. This evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
All evaluations were conducted by 14 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields. Each annotator was familiar with language model evaluation and prompt optimization procedures. The evaluation employed a 5-point Likert scale for reasoning quality (H1) and a 4-point scale for directional consistency (H2). The following subsections describe the evaluation setup and results.
5.3.1. Evaluating Multi-Agent Reasoning Quality (H1)
To assess whether MA-SAPO enhances reasoning quality, we randomly sampled 30 reasoning cases from the HelpSteer1 training set. Each case included two reasoning outputs: one generated by a single-agent baseline and one by MA-SAPO’s multi-agent reasoning pipeline, resulting in 60 reasoning outputs in total. These outputs were independently evaluated by 14 human annotators, yielding 840 individual evaluations (30 cases $\times$ 2 outputs $\times$ 14 annotators). Each output was rated along three qualitative dimensions:
  • Usefulness: How useful the response is in addressing the prompt.
  • Accuracy: How factually correct the reasoning content is.
  • Consistency: How coherent and logically structured the writing is.
Each dimension was rated on a scale from 1 (very low) to 5 (very high).
Results.
As summarized in Table 4, MA-SAPO outperformed the single-agent baseline across all three dimensions. The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements (*p* ¡ 0.05, paired $t$-test, $n=30$). These findings indicate that multi-agent reasoning yields more helpful, accurate, and consistent responses.
Table 4. Human evaluation of reasoning quality. MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency. Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
Method Usefulness Accuracy Consistency Mean
Single-Agent 3.64 3.63 3.81 3.69
MA-SAPO (Ours) 3.89 3.87 4.02 3.93
$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.23
5.3.2. Evaluating Directional Consistency of Optimized Prompts (H2)
To evaluate whether MA-SAPO preserves the semantic intent of prompts during optimization, we randomly sampled 40 prompt pairs, each consisting of an original prompt ($A$) and its optimized version ($B$) generated by MA-SAPO. Each pair was independently rated by 14 human annotators, resulting in a total of 560 evaluations (40 pairs $\times$ 14 annotators). Annotators rated the directional consistency of each pair on a 4-point Likert scale (1 = completely changed, 2 = partially changed, 3 = mostly preserved, 4 = fully preserved), indicating how well the optimized prompt maintained the intent, goal, and domain of the original.
Results.
Figure 3 visualizes the distribution of annotator ratings for directional consistency. Most scores are concentrated between 3 and 4, indicating that the optimized prompts largely preserved the original semantic intent. The mean rating was 3.36 with a standard deviation of 0.74 ($n=560$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability. The relatively moderate standard deviation further suggests that annotators consistently judged the optimized prompts as semantically aligned with their originals.
Figure 3. Distribution of directional consistency ratings (n=560). Ratings on a 4-point scale (1 = changed, 4 = preserved). The dashed line marks the mean (3.36) and the shaded band shows one standard deviation (SD = 0.74).
5.3.3. Summary of Human Evaluation
The human evaluation results highlight two complementary aspects of MA-SAPO. First, the multi-agent reasoning framework generates responses that are more useful, accurate, and consistent than those from a single-agent baseline, confirming the advantage of role-specialized reasoning. Second, MA-SAPO maintains strong semantic stability during prompt optimization, ensuring that improved clarity and structure do not distort the original task intent. Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in Section 5.3.
5.4. Cost and Latency
Table 5. Cost and Latency Comparison. MA-SAPO requires far fewer tokens and calls than other competitive multi-agent baselines, while maintaining superior optimization performance. Lower is better ($\downarrow$).
Method # Input Tokens $\downarrow$ # Output Tokens $\downarrow$ # Calls $\downarrow$
MAD 5,217 2,170 8
MARS 17,149 6,789 60
MA-SAPO (Ours) 4,968 526 2
To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks. Table 5 reports the average number of input tokens, output tokens, and calls per instance. We observe that existing multi-agent systems such as MAD and MARS incur substantial overhead. MAD requires around 5K input tokens, 2K output tokens, and 8 calls on average due to repeated multi-round debates. MARS is even heavier, consuming more than 17K input tokens, nearly 7K output tokens, and about 60 calls per instance, reflecting its iterative planner–critic–student workflow. In contrast, MA-SAPO is significantly more efficient: by constructing reasoning assets offline and adopting a single Analyzer to Refiner loop at test time, it reduces the runtime cost to roughly 5K input tokens, 0.5K output tokens, and only 2 calls per instance.
These results highlight that MA-SAPO not only improves optimization quality but also achieves a more favorable cost–latency trade-off. While MAD and MARS achieve moderate gains at the expense of heavy resource usage, MA-SAPO provides consistent improvements while keeping computational demands.
Figure 4. Case study illustrating MA-SAPO’s multi-agent prompt optimization process. Starting from an initial user prompt and evaluation scores, the framework retrieves reasoning assets, analyzes weaknesses, and iteratively refines the prompt. The regenerated response shows improved clarity, structure, and contextual depth compared to the original.
5.5. Case Study
To further illustrate the effectiveness and practical utility of MA-SAPO, we present a case study drawn from the HelpSteer2 validation dataset. [^7] Additional case studies demonstrating diverse application scenarios are available in our code repository. This example highlights MA-SAPO’s strengths in collaborative reasoning, context-aware prompt refinement, and robust adaptation across diverse input types, thereby providing concrete evidence of the framework’s ability to enhance prompt optimization in complex reasoning scenarios.
The case shown in Figure 4 demonstrates the end-to-end prompt optimization process. It begins with an initial input prompt, “What are 10 primary ways ChatGPT will change 4th–12th grade education in the United States?”, alongside the corresponding model response and evaluation scores. MA-SAPO then dynamically retrieves three semantically and thematically relevant exemplars from the training corpus (e.g., “applying ChatGPT in language learning and teaching”). These retrieved assets introduce diverse educational contexts and prompt formulations, implicitly guiding refinements such as narrowing task scope and aligning conceptual goals with actionable applications.
The Analyzer Agent identifies weaknesses in the original prompt notably, vague task specification and overly broad phrasing-both of which risk eliciting generic responses. Building on this analysis, the Refiner Agent reformulates the prompt to preserve its educational intent while incorporating a more actionable subtask, such as exploring business opportunities tied to ChatGPT-based assessment practices. The optimized prompt exhibits clearer structure, explicit segmentation, and richer contextual depth. When regenerated with the refined prompt, the model produces a response that is more coherent, pedagogically grounded, and practically insightful.
Overall, this case underscores how MA-SAPO integrates multi-agent reasoning and retrieved exemplars to achieve superior prompt optimization, outperforming conventional single-agent or static fine-tuning approaches in terms of coherence, interpretability, and alignment with user intent.
6. Conclusion
We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization. Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer to Refiner interactions. Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency. Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
Despite these advantages, MA-SAPO currently depends on the quality of human-annotated scores and a fixed semi-structured schema for asset construction, while also relying on a sparse BM25 retriever and a single reward model-factors that may overlook semantically relevant exemplars and introduce evaluator bias. In future work, we plan to improve robustness by integrating a dedicated Feedback Agent to validate asset quality and mitigate bias, as well as by developing hybrid dense–sparse retrieval strategies. We also aim to extend the framework to multi-turn prompts to enable more comprehensive evaluation.