### v2

# MA-SAPO: Multi-Agent Reasoning for Score-Aware Prompt Optimization

Abstract
Prompt optimization has become a practical way to improve the performance of Large Language Models (LLMs) without retraining. However, most existing frameworks treat evaluation as a black box, relying solely on outcome scores without explaining why prompts succeed or fail. Moreover, they involve repetitive trial-and-error refinements that remain implicit, offering limited interpretability or actionable guidance for systematic improvement. In this paper, we propose MA-SAPO: a new Multi-Agent Reasoning for Score Aware Prompt Optimization framework that links evaluation outcomes directly to targeted refinements. Specifically, in the Training Phase, multiple agents interpret evaluation scores, diagnose weaknesses, and generate concrete revision directives, which are stored as reusable reasoning assets. In the Test Phase, an analyzer agent retrieves relevant exemplars and assets for a new prompt, and a refiner agent applies evidence-based edits to improve the prompt and its response. By grounding optimization in structured reasoning, MA-SAPO ensures edits are interpretable, auditable, and controllable. Experiments on the HelpSteer1/2 benchmarks show that our framework consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent methods across multiple evaluation metrics.
1 Introduction
Large Language Models (LLMs) have emerged as powerful tools capable of tackling a wide range of tasks, from reasoning and summarization to complex dialogue and code generation (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). However, their performance remains highly sensitive to the wording and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Cao et al., 2024; Razavi et al., 2025). Consequently, prompt optimization has rapidly emerged as a practical alternative to costly retraining or parameter updates for improving model behavior.
Early research explored single-pass prompt optimization that effectively treats the LLM as its own optimizer. Approaches such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role conditioning via carefully crafted system instructions (Wu et al., 2023; Kong et al., 2024), and structured reasoning schemes including Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2024) demonstrated that encoding explicit reasoning structure or role signals in the prompt can materially improve downstream performance. Nevertheless, these methods are largely self-contained, relying on a single model instance to revise prompts without leveraging systematic feedback across iterations or external resources.
As LLMs have advanced, multi-agent frameworks have emerged to broaden viewpoints and coordinate specialized roles in prompt optimization (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025). By delegating subtasks such as decomposition, evaluation, and refinement to distinct agents, these systems extend beyond single-pass rewriting into more structured pipelines. In parallel, optimization-theoretic approaches (Li et al., 2025a; Cui et al., 2025) and label-efficient schemes (Xiang et al., 2025; Dong et al., 2025) aim to curb reliance on expensive feedback while scaling across tasks.
Despite these advances, several common limitations remain: (i) evaluation is often collapsed to scalar scores, obscuring why prompts succeed or fail; (ii) refinement remains largely trial-and-error, incurring high computational cost with limited transparency; (iii) reasoning, when used, is typically implicit rather than distilled into auditable artifacts that drive targeted edits; and (iv) pipelines prioritize score gains over interpretability and controllability, which hinders practitioners from analyzing trade-offs and making precise adjustments in real-world applications.
To address these limitations, we introduce MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly maps metric outcomes to actionable edits via structured reasoning artifacts. MA-SAPO has two phases. In the training phase, three agents convert annotated scores into reusable assets: (1) a Metric Explainer interprets evaluation dimensions, (2) a Diagnostician identifies error sources and trade-offs, and (3) an Action Synthesizer produces concrete edit directives. In the test phase, a retrieval-augmented pipeline applies these assets: an Analyzer contrasts the current prompt with retrieved exemplars to surface improvement opportunities, and a Refiner performs targeted edits grounded in diagnostic evidence, yielding interpretable and controllable optimization.
Experiments on HelpSteer1/2 (Wang et al., 2024c; b) show that MA-SAPO consistently surpasses single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, underscoring structured reasoning as an effective bridge between evaluation and optimization. Our contributions are:
  • A score-aware training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets for explanation, diagnosis, and edit directives.
  • MA-SAPO, a modular and compute-efficient framework that delivers interpretable, auditable, and controllable gains while reducing token and API-calls.
  • Extensive results on HelpSteer1/2 demonstrating consistent multi-metric improvements and outperforming six baselines in both effectiveness and efficiency.
2 Related work
2.1 Single-pass prompt optimization
Prompt optimization has become an effective way to improve LLM performance without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025b). Early work established prompts as key control variables for steering model behavior. In particular, few-shot prompting (Brown et al., 2020) showed that models can adapt to new tasks through in-context examples, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) improves factuality by injecting external knowledge. Together, these approaches laid the foundation for single-pass prompt optimization. Later work introduced reasoning-augmented prompting. Chain-of-Thought (CoT) (Wei et al., 2022) added intermediate reasoning steps, and Tree-of-Thought (ToT) (Yao et al., 2023) extended this into branching reasoning with self-evaluation. Other structured variants include Skeleton-of-Thought (SoT) (Ning et al., 2023), which separates planning from execution, and Graph-of-Thought (GoT) (Besta et al., 2024), which models reasoning as graph-structured trajectories. Evolutionary and heuristic methods have also been explored: PromptBreeder (Fernando et al., 2023) evolves prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines them via critique-and-synthesis loops.
2.2 Multi-agent based prompt optimization
Beyond single-pass methods, multi-agent systems have become an important paradigm for prompt optimization (Du et al., 2024; Liu et al., 2024; Chen et al., 2024; Li et al., 2024; Seo et al., 2026). Early agentic approaches such as Self-Ask (Press et al., 2023) introduced question decomposition and retrieval, while Multi-Agent Debate (MAD) (Du et al., 2024) improved reasoning through iterative answer exchange and critique. Reflection-based methods (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by incorporating feedback from prior attempts. Building on these ideas, later frameworks structured agent roles and workflows more explicitly (Li et al., 2023) and began targeting prompt optimization itself. For example, MASS (Zhou et al., 2025) searches over agent topologies with local and global prompt refinement, MARS (Zhang et al., 2025) uses Planner and Teacher–Critic–Student interactions for iterative revision, and MAPGD (Han et al., 2025) combines pseudo-gradient feedback with bandit-based exploration. However, many existing methods still treat evaluation as a black box and rely on costly trial-and-error refinement. MA-SAPO instead grounds optimization in score-aware reasoning assets, making revisions more interpretable, auditable, and actionable.
3 Methodology
Figure 1: MA-SAPO pipeline overview. Three training agents convert annotated prompt–response pairs into reusable Reasoning Assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$. At test time, retrieved assets guide the Analyzer $G_{\mathrm{ana}}$ and Refiner $G_{\mathrm{ref}}$ to produce an optimized prompt $\hat{p}$ with higher scores across all five evaluation dimensions.
3.1 Preliminaries
Let the training dataset be denoted as
$\mathcal{D}_{train}=\{\tau_{i}\}_{i=1}^{N},\quad\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i}),$
where each entry $\tau_{i}$ consists of a prompt $p_{i}$, its associated response $r_{i}$, and a set of scores $\mathcal{S}_{i}=\{s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\}$, corresponding to the dimensions of helpfulness, correctness, coherence, complexity, and verbosity. The test set is denoted as
$\mathcal{D}_{test}=\{p_{j}\}_{j=1}^{M},$
where only prompts are available, and corresponding responses are generated during evaluation. Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can be retrieved to optimize new prompts and their corresponding responses. Thus, the training phase serves as data generation, while the test phase operates as a retrieval-augmented optimization pipeline.
Throughout this framework, we employ several specialized agents denoted as $G$. Each agent is designed with a distinct functional role: $G_{\text{exp}}$ explains metric-level outcomes, $G_{\text{diag}}$ diagnoses weaknesses and trade-offs, $G_{\text{syn}}$ synthesizes actionable directives, $G_{\text{ana}}$ analyzes retrieved examples, and $G_{\text{ref}}$ refines prompts into optimized versions. These agents are executed in sequential order within their respective phases, and their outputs are explicitly stored for later retrieval and reasoning. The overview of the MA-SAPO framework is shown in Figure 1.
3.2 Training phase: reasoning asset construction
In the training phase, three agents are executed sequentially for each $\tau_{i}\in\mathcal{D}_{train}$. All three agents operate on the same triplet $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$, ensuring that their reasoning is consistently grounded in both the input and output of the original instance.
3.2.1 Metric explainer agent $G_{\text{exp}}$
The Metric Explainer Agent generates natural language justifications for the assigned scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
$\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$ (1)
where $P_{\text{exp}}$ is the system prompt guiding explanation. Each card explicitly outlines the reason why $\tau_{i}$ received its annotated scores. The reasoning card $\mathcal{C}_{i}$ then serves as input to the next agent, together with $\tau_{i}$.
3.2.2 Diagnostician agent $G_{\text{diag}}$
The Diagnostician Agent operates on $\tau_{i}$ and the reasoning card $\mathcal{C}_{i}$. It extends $\mathcal{C}_{i}$ by analyzing metric-level weaknesses and trade-offs. $G_{\text{diag}}$ produces a diagnostic summary $\mathcal{D}_{i}$ that identifies: (1) the key causes of low-scoring dimensions, and (2) trade-offs across metrics (e.g., verbosity vs. coherence).
$\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$ (2)
The diagnostic summary $\mathcal{D}_{i}$ is then passed to the next stage, along with $\tau_{i}$.
3.2.3 Action synthesizer agent $G_{\text{syn}}$
The Action Synthesizer Agent receives $\tau_{i}$ together with $\mathcal{C}_{i}$ and $\mathcal{D}_{i}$. It converts these enriched insights into actionable edit directives (EDs). Each directive corresponds to a concrete modification strategy for improving $\tau_{i}$:
$\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$ (3)
where $\mathcal{E}_{i}=\{e_{1},e_{2},\dots,e_{m}\}$ is a set of recommended edits.
Definition (reasoning assets).
For notational convenience, we define the collection of training outputs for instance $i$ as:
$\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}).$
As a result, $\mathcal{R}_{i}$ is stored as reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase. Importantly, these reasoning assets are designed as semi-structured text, which makes them machine-parseable and directly usable by downstream agents in the test phase. For instance, edit directives often follow a pattern such as “add a constraint to …”, which the refiner agent can reliably identify and execute.
3.3 Test phase: retrieval and reformulation
In the test phase, optimization is performed via retrieval-augmented generation. Given a new prompt $p_{\text{test}}$, we adopt a sparse lexical retriever that ranks training prompts based on term-frequency and inverse document-frequency statistics. Each training prompt $p_{j}$ is represented in the sparse lexical space, and a relevance score is computed between $p_{\text{test}}$ and $p_{j}$ according to token overlap weighted by discriminative importance. The top-$k$ most relevant prompt-response pairs are then retrieved along with their corresponding reasoning assets:
$\{(\tau_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
3.3.1 Analyzer agent $G_{\text{ana}}$
The Analyzer Agent compares $p_{\text{test}}$ against retrieved examples to identify improvement points. It outputs an improvement report $\mathcal{A}$:
$\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{\text{ana}}).$ (4)
The Analyzer Agent transforms raw retrieval into structured insights rather than merely transferring edits from similar prompts. The agent highlights concrete weaknesses and improvement opportunities by contrasting the test prompt with retrieved reasoning assets.
3.3.2 Refiner agent $G_{\text{ref}}$
The Refiner Agent regenerates an optimized prompt $\hat{p}$ by incorporating the improvement report:
$\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$ (5)
It operationalizes the analyzer’s insights into a concrete optimized prompt. Compared to a direct rewriting approach, which may introduce irrelevant or unjustified changes, the refiner explicitly conditions on the improvement report, producing refinements that are focused, justifiable, and consistent with diagnostic evidence. The final optimized prompt $\hat{p}$ is then used to generate a corresponding optimized response $\hat{r}$. [^1] The detailed prompts and specific role descriptions for agents are provided in Appendix H.
4 Experimental design and setup
4.1 Datasets
We conduct experiments on the HelpSteer datasets (Wang et al., 2024c; b), which provide human-annotated prompt-response pairs scored on five quality dimensions using a 0–4 scale. [^2] Detailed descriptions of the five annotation dimensions are provided in Table 4 in Appendix A. Both datasets follow the same annotation schema but differ in size and construction, as summarized in Table 1. We build the retrieval corpus solely from the HelpSteer2 training set, which offers higher-quality annotations, broader prompt diversity, and preference labels between responses that provide richer supervision for our Analyzer and Refiner agents. HelpSteer2 also serves as a standard benchmark for reward modeling. Evaluation is conducted on the validation splits of HelpSteer1 and HelpSteer2.
Dataset Train Val Multi-turn Preference Ann.
HelpSteer1 35.3k 1.79k × ×
HelpSteer2 20.3k 1.04k
Table 1: Statistics of the HelpSteer1/2 datasets. Both share the same annotation schema of five quality dimensions scored on a 0–4 scale.
4.2 Models used
Reasoning–asset generation (corpus construction).
To construct the retrieval corpus, we employ the three reasoning agents ($G_{\text{exp}}$, $G_{\text{diag}}$, $G_{\text{syn}}$) using the o4-mini reasoning model from OpenAI (Achiam et al., 2023). This model is used exclusively for reasoning asset construction rather than downstream inference. All generations follow default API hyperparameters. For each training instance $\tau_{i}$, it produces a reasoning card $\mathcal{C}_{i}$, a diagnostic summary $\mathcal{D}_{i}$, and a set of edit directives $\mathcal{E}_{i}$, which together form the stored reasoning assets $\mathcal{R}_{i}$.
Generation models for evaluation.
For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model Llama3-8B-Instruct (Dubey et al., 2024) as the backbone models. These models are responsible for generating (i) optimized prompts $\hat{p}$ from the Refiner Agent and (ii) the corresponding optimized responses $\hat{r}$ conditioned on $\hat{p}$. Both models are configured with temperature =0=0 to ensure deterministic decoding, and all other hyperparameters remain at their defaults.
Judge model.
We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024a) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality. For any prompt-response pair, the judge model returns five HelpSteer-aligned scores: [^3] A detailed description of the ArmoRM model architecture, multi-objective training methodology, and alignment with HelpSteer dimensions is provided in Appendix C.
$\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\},\quad s_{m}\in[0,4].$
Retrieval.
We adopt a sparse retriever based on BM25 (Robertson et al., 2009) over prompt text. For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^4] A detailed analysis of $k$ values and their effects on performance is provided in Section 5.2.1. and attach their paired responses and reasoning assets:
$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k=3}.$ (6)
4.3 Evaluation metrics
We evaluate each candidate pair $(p,r)$ on the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$. Scores are normalized as $\tilde{s}_{m}=s_{m}/4\in[0,1]$, and the overall quality score is computed as the average across all dimensions:
$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$ (7)
This yields a single composite score in $[0,1]$ representing overall response quality. We also conduct a human evaluation with 17 expert annotators to assess (H1) the usefulness, accuracy, and consistency of the reasoning, and (H2) whether optimized prompts preserve the original intent. [^5] Human evaluation results are reported in Section 5.3, with detailed protocols and per-dimension scores in Appendix F.
4.4 Baselines
We compare our method with six baselines in three categories: single-pass prompt optimization without retrieval (Direct Generation, CoT (Wei et al., 2022), Role Assignment (Wu et al., 2023)), retrieval-augmented optimization (RAG (Lewis et al., 2020)), and multi-agent optimization (MAD (Du et al., 2024), MARS (Zhang et al., 2025)). For all methods, the model first generates an optimized prompt and then produces the final response using the same backbone LLM. We evaluate all baselines with GPT-4o and Llama3-8B-Instruct, using temperature 0 and default decoding settings. For RAG, we retrieve the top-$10$ training examples with BM25; for MAD, we use $2$ agents and $3$ debate rounds; and for MARS, we use 3344 planning steps and $5$ refinement iterations. [^6] Baseline details are provided in Appendix B.
5 Experimental results and analysis
5.1 Main results
Methods GPT-4o Llama3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1
Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
RAG (sparse, k=10k{=}10) (2020) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
Chain-of-Thought (CoT) (2022) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
Role Assignment (2023) 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
MAD (2024) 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
MARS (2025) 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
MA-SAPO (Ours) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2
Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
RAG (sparse, k=10k{=}10) (2020) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
Chain-of-Thought (CoT) (2022) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
Role Assignment (2023) 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
MAD (2024) 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
MARS (2025) 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
MA-SAPO (Ours) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
Table 2: Main results on HelpSteer1 and HelpSteer2. Columns report the five HelpSteer metrics (Help, Corr, Coh, Comp, Verb) normalized to $[0,1]$ and their mean (Avg). Methods are grouped into (i) single-pass prompting, (ii) retrieval-augmented generation (no reasoning assets), and (iii) multi-agent frameworks. For each dataset and backbone block, the best value in each metric column is shown in bold, and the second-best is underlined.
In our main experiments, we compare MA-SAPO with three groups of baselines: (1) single-pass prompting methods (Direct Generation, CoT, and Role Assignment), (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks (MAD and MARS). As shown in Table 2, MA-SAPO consistently achieves the best results across both HelpSteer1 and HelpSteer2 under both GPT-4o and Llama3-8B. Single-pass methods remain limited because they revise prompts without diagnostic grounding, often leading to brittle or inconsistent improvements. Role Assignment provides more explicit guidance, but still depends heavily on predefined roles. RAG benefits from retrieved exemplars, but without explicit reasoning about their usefulness, its refinements remain heuristic and less reliable.
Figure 2: KDE score distributions of Direct Generation and MA-SAPO across five HelpSteer dimensions and their average, aggregated over GPT-4o and Llama3-8B. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution. MA-SAPO consistently shifts the score distribution toward higher values.
Methods GPT-4o Llama3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1
$k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
$k=2$ 0.5211 0.6287 0.8591 0.4960 0.7318 0.6473 0.4002 0.4747 0.8236 0.4726 0.8384 0.6019
$k=4$ 0.5204 0.6266 0.8606 0.4965 0.7369 0.6482 0.4028 0.4778 0.8272 0.4708 0.8389 0.6035
Test Agents Combination 0.4672 0.5415 0.8619 0.3670 0.6403 0.5756 0.4649 0.5203 0.8323 0.4363 0.7249 0.5957
MA-SAPO (kk=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2
$k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
$k=2$ 0.5107 0.6044 0.8504 0.5158 0.7450 0.6452 0.3980 0.4713 0.8268 0.4845 0.8396 0.6041
$k=4$ 0.5066 0.6012 0.8480 0.5202 0.7451 0.6442 0.3947 0.4685 0.8251 0.4786 0.8395 0.6013
Test Agents Combination 0.5404 0.6012 0.8891 0.4455 0.7398 0.6432 0.4379 0.4720 0.7968 0.4918 0.7321 0.5861
MA-SAPO (kk=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
Table 3: Ablation results over agent number $k$ and agent-combination testing on HelpSteer1 and HelpSteer2. Best scores are shown in bold and second-best scores are underlined.
Multi-agent baselines such as MAD and MARS achieve stronger performance by coordinating multiple roles, but they still rely on repeated interaction cycles and treat evaluation mainly as a scalar target. In contrast, MA-SAPO converts evaluation outcomes into reusable reasoning assets and uses them to support evidence-backed refinement through the Analyzer-to-Refiner pipeline, yielding stronger and more interpretable improvements. For example, on HelpSteer1 with GPT-4o, MA-SAPO achieves an average score of 0.6486, outperforming MARS (0.5096) and RAG (0.4784). Similar trends hold across HelpSteer2 and Llama3-8B, and Figure 2 further shows that these gains reflect a consistent shift toward higher scores across dimensions rather than a few outlier cases. [^7] Per-backbone KDE distributions are provided in Appendix E. This suggests that MA-SAPO improves multiple quality aspects simultaneously, rather than over-optimizing a single metric.
5.2 Ablation study
5.2.1 Top-k retrieval variations
To further analyze the robustness of MA-SAPO, we examine two key design choices: retrieval depth and test-phase agent composition. First, to study how the number of retrieved training examples affects performance, we vary the retrieval depth $k\in\{1,2,3,4\}$. Table 3 reports results for both GPT-4o and Llama3-8B across HelpSteer1 and HelpSteer2. Across both backbones, increasing $k$ from 1 to 3 generally improves performance, with $k=3$ achieving the best average balance across the five metrics. When $k=1$, the Analyzer is limited to a single exemplar, restricting the diversity of available reasoning assets. Increasing $k$ broadens the evidence base, but $k=4$ shows slight declines in some dimensions, suggesting that retrieving too many exemplars can introduce noise and conflicting signals. Overall, these results indicate a clear trade-off between contextual diversity and diagnostic noise, with $k=3$ providing the most reliable setting in our experiments.
5.2.2 Test agents combination
We subsequently assess the effect of test-phase agent modularity by merging the two agents into a single combined agent, denoted as Test Agents Combination. As shown in Table 3, this variant remains competitive for both GPT-4o and Llama3-8B on HelpSteer1 and HelpSteer2, and occasionally surpasses the full framework on individual metrics. However, its average performance consistently falls below that of the full MA-SAPO, indicating that it cannot reliably reproduce the benefits of the structured Analyzer-to-Refiner workflow. By separating evidence-grounded diagnosis from targeted refinement, the modular design supports a more interpretable and controllable process and leads to more consistent gains across datasets and model backbones.
5.3 Qualitative analysis
We additionally conducted a human evaluation to examine two complementary aspects of MA-SAPO. [^8] The full evaluation protocol, per-dimension scores, and the rating distribution are provided in Appendix F. A total of 17 annotators, comprising data scientists, AI engineers, and graduate students in data science or related fields, participated in the study. For reasoning quality (H1), MA-SAPO achieved a mean score of 3.93 on a 5-point scale, outperforming the single-agent baseline by 0.24 points, with notable gains in usefulness and accuracy. For directional consistency (H2), annotators assessed whether optimized prompts retained the original intent, yielding a mean rating of 3.36 out of 4.0, indicating that prompt optimization largely preserves semantic alignment. Taken together, the human evaluation supports the quantitative findings and demonstrates that MA-SAPO improves output quality without distorting the original task intent.
5.4 Cost and efficiency analysis
Figure 3: Efficiency–performance trade-off of MAD, MARS, and MA-SAPO. Panel (a) plots average input tokens per instance against average score, and panel (b) plots execution time per instance against average score, under GPT-4o and Llama3-8B. Averaged over HelpSteer1 and HelpSteer2, MA-SAPO delivers the highest scores with the lowest token cost and substantially lower latency than MARS.
To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks. Figure 3(a) reports the average input tokens per instance against average score. MAD incurs moderate overhead but yields relatively weak overall performance, while MARS improves performance at a substantially higher token cost, consuming over 17K input tokens per instance. In contrast, MA-SAPO reduces token consumption to approximately 5K input tokens per instance by shifting reasoning asset construction offline and using a single Analyzer-to-Refiner loop at test time, while achieving the highest average scores under both GPT-4o and Llama3-8B. Figure 3(b) further illustrates the trade-off between execution time per instance and average score. MAD achieves the shortest execution time at approximately 5 seconds per instance, yet delivers the lowest average performance among the three methods. MARS incurs the highest latency, requiring roughly 85 to 90 seconds per instance. By contrast, MA-SAPO maintains an execution time of approximately 40 seconds per instance, which is less than half that of MARS, while delivering substantially higher scores. Taken together, these results show that MA-SAPO achieves the most favorable efficiency-performance balance across both token cost and latency. [^9] Detailed model inference hardware specifications are provided in Appendix D.
6 Conclusion
We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization. Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer-to-Refiner interactions. [^10] Practical implications are discussed in Appendix G. Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency. Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
Limitations
Although MA-SAPO achieves superior performance compared to a range of single-agent reasoning baselines and interaction-based agent frameworks, these results do not imply that MA-SAPO uniformly dominates all baselines across tasks and conditions. MA-SAPO leverages retrieved exemplars as evidence for diagnosis and prompt rewriting. However, retrieval may suffer from limited recall under paraphrase or lexical mismatch, causing the system to rely on less relevant assets and creating a bottleneck at the retrieval stage. Moreover, the single-path, sequential multi-agent pipeline is prone to cascading errors and provides limited exploration, since it does not explicitly branch over, compare, and select among alternative diagnostic/edit trajectories. Future work could improve robustness by adding verifier-based artifact validation, uncertainty-aware gated updates, and constrained rewriting, as well as by introducing multi-path candidate generation with iterative verification.
Appendix
Appendix A Datasets detail
We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2024c; b), which provide human-annotated prompt-response pairs for evaluating prompt optimization in large language models. Both datasets share the same annotation schema across five quality dimensions, but differ in size, design, and annotation depth, as summarized in Table 4. We build our retrieval corpus from the HelpSteer2 training set and conduct evaluation on the validation splits of both datasets.
Attribute Description
Helpfulness Overall helpfulness of the response to the prompt.
Correctness Inclusion of all pertinent facts without errors.
Coherence Consistency and clarity of expression.
Complexity Intellectual depth required to write the response.
Verbosity Amount of detail included in the response relative to what is asked.
Table 4: Overview of HelpSteer1/2 annotation dimensions. Each scored on a 0–4 scale.
Figure 4: Score distributions across five annotation dimensions for HelpSteer1 (left) and HelpSteer2 (right). Each bar shows the proportion of samples assigned each score from 0 to 4. Coherence is heavily concentrated at higher scores in both datasets, while complexity and verbosity exhibit broader distributions across the full range.
A.1 HelpSteer1
HelpSteer1 contains 35.3k training and 1.79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0–4 scale. Prompts cover a broad range of task types including rewriting, summarization, classification, extraction, and question answering, with up to four responses generated per prompt. Each response was independently reviewed by multiple annotators, providing reliable quality estimates across diverse prompt-response styles.
A.2 HelpSteer2
HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024b). Compared to HelpSteer1, it employs a substantially larger annotator pool and a more diverse set of response sources, with each sample receiving a minimum of three independent annotations. We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field. As shown in Figure 4, HelpSteer2 exhibits more concentrated distributions toward higher scores, particularly in coherence, reflecting the higher quality of responses.
Appendix B Baseline details
We compare our method against six baselines from three categories: single-pass prompting, retrieval-augmented generation, and multi-agent frameworks. For all methods, the model first produces an optimized prompt from the initial prompt, and the final response is then generated from the optimized prompt using the same backbone LLM.
B.1 Single-pass prompting
This category includes Direct Generation, Chain-of-Thought (CoT) (Wei et al., 2022), and Role Assignment (Shanahan et al., 2023). These methods optimize the prompt without retrieving external examples. Direct Generation rewrites the input prompt into an optimized prompt in a single pass. Chain-of-Thought (CoT) generates intermediate reasoning steps before producing the optimized prompt. Role Assignment prepends a predefined role or persona to the input prompt and generates the optimized prompt in a single pass.
B.2 Retrieval-augmented generation
This category includes RAG (Lewis et al., 2020). The model is provided with the top-$k=10$ retrieved training examples as references when generating the optimized prompt. No additional reasoning assets are used.
B.3 Multi-agent framework
This category includes MAD (Du et al., 2024) and MARS (Zhang et al., 2025). MAD uses multiple agents to propose candidate prompt revisions, exchange feedback over several rounds, and select the final optimized prompt. MARS uses a hierarchical multi-agent framework that first plans the optimization procedure and then iteratively refines the prompt.
For all baselines, we use GPT-4o and Llama3-8B-Instruct as the backbone models. We set the decoding temperature to 0 and keep all other generation settings at their default values. For MAD, we use $2$ agents and $3$ debate rounds. For MARS, the planner produces $3$ to $4$ planning steps, followed by $5$ refinement iterations.
Appendix C LLM-as-a-judge
To automatically evaluate the quality of responses generated from our HelpSteer-based prompt generation framework, we use ArmoRM-Llama3-8B-v0.1 as the LLM-as-a-judge. ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) is a Llama-3 8B-based reward model that departs from conventional scalar preference modeling by decomposing response quality into multiple interpretable reward objectives. Specifically, the model first learns attribute-level rewards from multi-dimensional absolute-rating data through an Absolute-Rating Multi-Objective Reward Model (ArmoRM), and then aggregates them into a final preference score using a Mixture-of-Experts (MoE) gating network conditioned on the prompt context.
This model is particularly suitable for our evaluation setting because it directly incorporates the five HelpSteer annotation dimensions (helpfulness, correctness, coherence, complexity, and verbosity) as internal reward objectives. Since the model is trained to predict these HelpSteer-related attributes as part of its multi-objective reward formulation, it is well aligned with our task of evaluating prompt-response pairs derived from a HelpSteer-based framework. In addition, its two-stage design separates attribute-level reward prediction from final scalar aggregation, which provides a more interpretable evaluation process than conventional single-score preference models.
Empirically, ArmoRM-Llama3-8B-v0.1 achieves an overall score of 89.0 on RewardBench, outperforming the corresponding Llama-3 8B Bradley–Terry reward model (83.6). These reported results support our use of ArmoRM-Llama3-8B-v0.1 not only as an open and efficient judge model, but also as a reward model whose internal objectives are structurally aligned with the HelpSteer evaluation dimensions.
Appendix D Hardware Specifications
Model Hardware Power (TDP) Access
Meta-Llama-3-8B-Instruct NVIDIA H100-80GB 700W Local GPU
GPT-4o OpenAI API
Table 5: Hardware specifications used for model inference.
Table 5 summarizes the hardware configurations used for model inference in our experiments. For the open-source model, inference with Meta-Llama-3-8B-Instruct was conducted on NVIDIA H100-80GB GPU. For the proprietary model, GPT-4o was accessed via the OpenAI API, and no local GPU resources were required.
Appendix E Per-backbone score distributions
Figures 5 and 6 decompose the aggregated KDE distributions from Figure 2 by backbone model, reporting score distributions for GPT-4o and Llama3-8B separately. In each figure, the five HelpSteer dimensions and their average are shown as individual subplots, comparing Direct Generation against MA-SAPO across both HelpSteer1 and HelpSteer2.
E.1 GPT-4o
Figure 5: KDE score distributions for GPT-4o. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution.
Under GPT-4o, MA-SAPO produces pronounced rightward shifts across all five dimensions. Helpfulness and complexity show the largest gains, with Direct Generation distributions concentrated around scores of 1.2–1.3 and MA-SAPO shifting the mass to 2.1–2.2. Verbosity exhibits the most dramatic absolute shift, with Direct Generation peaking near 1.9 and MA-SAPO reaching approximately 3.0. Correctness also improves substantially, with the distribution peak moving from roughly 1.8 to 2.5 while becoming more tightly concentrated. Coherence is the exception, as Direct Generation already achieves high scores near 3.0, leaving limited headroom for further improvement. Nevertheless, MA-SAPO still achieves a visible upward shift in the distribution mean. The Average subplot confirms that these per-metric gains compound into a consistent overall improvement across both HelpSteer1 and HelpSteer2.
E.2 Llama3-8B
Figure 6: KDE score distributions for Llama3-8B. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution.
Under Llama3-8B, the directional pattern is consistent with GPT-4o, though with notable differences in distribution shape. Correctness under Direct Generation exhibits a broad, flat distribution across low scores, whereas MA-SAPO produces a sharp, concentrated peak near 2.0, indicating not only higher scores but also more consistent output quality. Verbosity shows the most dramatic shift among all metrics, with the MA-SAPO distribution peak reaching approximately 3.4 compared to roughly 1.8 for Direct Generation. Helpfulness and complexity also shift rightward, though with somewhat smaller absolute gains than under GPT-4o. As with GPT-4o, coherence shows the most modest improvement, consistent with its already high baseline scores.
E.3 Cross-backbone comparison
Across both backbones, MA-SAPO consistently shifts score distributions toward higher values in all five dimensions. Verbosity, complexity, helpfulness, and correctness show the largest shifts, while coherence improves more modestly due to the high baseline scores already achieved by Direct Generation. One notable difference between backbones is that Llama3-8B shows a sharper distributional improvement in correctness, with Direct Generation producing a much flatter spread compared to GPT-4o. Overall, the per-backbone breakdown confirms that the improvements observed in the aggregated figure are consistent across model families.
Appendix F Qualitative analysis details
This appendix provides the full details of the human evaluation summarized in Section 5.3. The evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
All evaluations were conducted by 17 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields. Each annotator was familiar with language model evaluation and prompt optimization procedures. The evaluation employed a 5-point Likert scale for reasoning quality (H1) and a 4-point scale for directional consistency (H2).
F.1 Evaluating multi-agent reasoning quality (H1)
To assess whether MA-SAPO enhances reasoning quality, we randomly sampled 30 reasoning cases from the HelpSteer1 training set. Each case included two reasoning outputs: one generated by a single-agent baseline and one by MA-SAPO’s multi-agent reasoning pipeline, resulting in 60 reasoning outputs in total. These outputs were independently evaluated by 17 human annotators, yielding 1020 individual evaluations (30 cases $\times$ 2 outputs $\times$ 17 annotators). Each output was rated along three qualitative dimensions:
  • Usefulness: How useful the response is in addressing the prompt.
  • Accuracy: How factually correct the reasoning content is.
  • Consistency: How coherent and logically structured the writing is.
Each dimension was rated on a scale from 1 (very low) to 5 (very high).
Results.
As summarized in Table 6, MA-SAPO outperformed the single-agent baseline across all three dimensions. The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements ($p<0.05$, paired $t$-test, $n=30$). These findings indicate that multi-agent reasoning yields more helpful, accurate, and consistent responses.
Method Usefulness Accuracy Consistency Mean
Single-Agent 3.64 3.63 3.81 3.69
MA-SAPO (Ours) 3.89 3.87 4.02 3.93
$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.24
Table 6: Human evaluation of reasoning quality. MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency. Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
F.2 Evaluating directional consistency of optimized prompts (H2)
To evaluate whether MA-SAPO preserves the semantic intent of prompts during optimization, we randomly sampled 40 prompt pairs, each consisting of an original prompt ($A$) and its optimized version ($B$) generated by MA-SAPO. Each pair was independently rated by 17 human annotators, resulting in a total of 680 evaluations (40 pairs $\times$ 17 annotators). Annotators rated the directional consistency of each pair on a 4-point Likert scale (1 = completely changed, 2 = partially changed, 3 = mostly preserved, 4 = fully preserved), indicating how well the optimized prompt maintained the intent, goal, and domain of the original.
Results.
Figure 7 visualizes the distribution of annotator ratings for directional consistency. Most scores are concentrated between 3 and 4, indicating that the optimized prompts largely preserved the original semantic intent. The mean rating was 3.36 with a standard deviation of 0.74 ($n=680$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability. The relatively moderate standard deviation further suggests that annotators consistently judged the optimized prompts as semantically aligned with their originals.
Figure 7: Distribution of directional consistency ratings (n=680). Ratings on a 4-point scale (1 = changed, 4 = preserved). The dashed line marks the mean (3.36) and the shaded band shows one standard deviation (SD = 0.74).
F.3 Summary of human evaluation
The human evaluation results highlight two complementary aspects of MA-SAPO. First, the multi-agent reasoning framework generates responses that are more useful, accurate, and consistent than those from a single-agent baseline, confirming the advantage of role-specialized reasoning. Second, MA-SAPO maintains strong semantic stability during prompt optimization, ensuring that improved clarity and structure do not distort the original task intent. Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in the main text.
Appendix G Implications
This work shows that prompt optimization can move beyond black-box score chasing by explicitly converting evaluation outcomes into reusable reasoning assets (explanations, diagnoses, and edit directives) that support targeted, evidence-backed edits. In practice, MA-SAPO offers a more interpretable and controllable workflow: rather than relying on trial-and-error rewriting, practitioners can trace changes to retrieved exemplars and their metric-linked rationales, which is valuable when prompt updates must be reviewed, audited, or kept consistent across versions. Moreover, MA-SAPO enables cost-aware scaling by shifting most computation offline: once the reasoning-asset corpus is built, new prompts can be optimized with lightweight retrieval plus an Analyzer-to-Refiner step, making it suitable for maintaining prompt libraries in real systems (e.g., assistants, support bots, enterprise applications) while reducing expensive multi-round agent interactions. For instance, the key contribution of MA-SAPO is a contrastive mechanism that compares, verifies, and selects among multiple LLM-generated candidates, rather than relying on advanced retrieval. As a result, even with BM25 alone, our framework remains competitive while being cheaper and faster to run. Consequently, MA-SAPO turns evaluation traces into retrievable, auditable reasoning assets, enabling repeatable prompt updates with predictable cost.
Appendix H Prompt template & design
In this section, we provide the complete prompt templates used to orchestrate the collaborative reasoning of each agent within the MA-SAPO framework. As briefly discussed in the main text, these specialized instructions are categorized into two distinct phases. The training phase prompts guide the Metric Explainer, Diagnostician, and Action Synthesizer agents in generating structured reasoning assets from evaluation scores. Conversely, the test phase prompts direct the Analyzer and Refiner agents to sequentially optimize the user’s initial prompt based on the retrieved assets. The detailed role descriptions, core tasks, and strict output constraints for each agent are presented below.
H.1 Training phase prompts
Prompt template: Metric explainer agent
Role: You are an evaluation explainer. Task: Explain WHY each of the five scores (0–4) was assigned to the model response. Provide a cohesive paragraph that bridges the gap between numerical scores and actionable improvements. Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} Strict Rules: Output must be a single paragraph in plain text. Cover metrics in order: helpfulness $\rightarrow$ correctness $\rightarrow$ coherence $\rightarrow$ complexity $\rightarrow$ verbosity. Ground every rationale specifically in the provided response evidence. Maintain a neutral and specific tone.
Table 7: Prompt template for the metric explainer agent.
Prompt template: Diagnostician agent
Role: You are a precise evaluation doctor for LLM outputs. Task: Diagnose the primary strengths and weaknesses of the response based on metric explanations. Prescribe concrete fixes to resolve identified reasoning or alignment gaps. Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} [REASONING_PARAGRAPH]: {reasoning_paragraph} Strict Rules: Rely strictly on the prompt, response, and provided reasoning. Address each metric once with a brief, specific rationale. Flag any mismatches between the numerical scores and the qualitative reasoning. Conclude with a crisp prescription for the required response elements.
Table 8: Prompt template for the diagnostician agent.
Prompt template: Action synthesizer agent
Role: You are a suggestion provider. Task: Synthesize the diagnosis into a single actionable prompt suggestion. Focus on enhancing the prompt’s structural and contextual depth without altering the original intent. Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} [DIAGNOSIS]: {diagnosis_text} Strict Rules: Output ONLY the improvement suggestion (no conversational filler). Maintain the core topic and original objective of the user prompt. Focus on actionable directions (e.g., adding constraints, requesting specific formats). Balance the five metrics to guide the model toward a higher-quality response.
Table 9: Prompt template for the action synthesizer agent.
H.2 Test phase prompts
Prompt template: Analyzer agent
Role: You are the analyzer agent. Task: Conduct an independent evaluation of the prompt’s inherent qualities. Integrate insights from the three previous agents (Metric Explainer, Diagnostician, Action Synthesizer) into a unified analysis. Inputs: [USER_PROMPT]: {prompt} [RETRIEVED_CONTENT]: {retrieved_content} (Aggregated agent outputs) Strict Rules: Prioritize the Action Synthesizer’s output when synthesizing final improvement points. Eliminate redundant observations across agent outputs. Ensure all proposed improvements are concrete and actionable. Output the final analysis in a structured, cohesive paragraph.
Table 10: Prompt template for the analyzer agent.
Prompt template: Refiner agent
Role: You are the refiner agent. Task: Optimize the original prompt based on the Analyzer Agent’s feedback. Produce a final, high-performance prompt that maximizes clarity and task effectiveness. Inputs: [USER_PROMPT]: {prompt} [ANALYZER_FEEDBACK]: {analyzed_result} Strict Rules: Directly incorporate the prioritized improvements from the feedback. Preserve the original intent and task objective perfectly. Remove any unnecessary elaboration to ensure the prompt remains concise. Output ONLY the final optimized prompt.
Table 11: Prompt template for the refiner agent.
Appendix I Case study
To further illustrate the effectiveness and practical utility of MA-SAPO, we present three case studies drawn from the HelpSteer2 validation dataset. [^11] Additional case studies demonstrating diverse application scenarios are available in our code repository. Rather than merely reiterating the optimization pipeline, these examples highlight MA-SAPO’s versatile problem-solving capabilities across different types of prompt deficiencies: (1) structuring broadly scoped queries, (2) contextualizing highly open-ended requests, and (3) disambiguating subjective terminology. Across all scenarios, the core multi-agent pipeline remains consistent: specialized agents (e.g., the Analyzer and Refiner) collaboratively evaluate retrieved exemplars to dynamically synthesize a targeted optimization strategy. By examining these distinct scenarios, we provide concrete evidence of the framework’s ability to guide language models toward more coherent, interpretable, and intent-aligned responses.
I.1 Case 1: structuring broadly scoped tasks
Figure 8: Case 1. For a broad query on ChatGPT’s educational impact, retrieved exemplars signal how to frame actionable tasks. The system identifies missing operational boundaries and restructures the prompt to demand a specific format (a 10-point list with explanations). This structural constraint shifts the final response from generic statements to pedagogically grounded, practical insights.
I.2 Case 2: contextualizing open-ended requests
Figure 9: Case 2. For an under-specified request for “new novel ideas,” retrieved exemplars highlight the need for a thematic anchor. The framework overhauls the prompt by enforcing explicit domain constraints and multi-perspective exploration. This enriched context transforms the final response from boilerplate advice into deep, domain-specific conceptualization.
I.3 Case 3: disambiguating subjective terminology
Figure 10: Case 3. For a query asking for the “strongest” character, retrieved exemplars provide cues for multi-dimensional evaluation. The framework refines the prompt by decomposing the subjective term “strongest” into explicit criteria (physical, strategic, and magical abilities). These multi-faceted constraints yield a final answer that replaces a simple name-drop with a nuanced, criteria-driven assessment.